An acronym for Zion Occupied Government. Probably stands for some other things, too.
Whats the mean of a set of data and explain how to find it?
take the sum of all the numbers ∑x
and divide it by the number of numbers in the data set. nor ∑x/n
A matched test is appropriate only if the two variables are measures taken for the same subject.
How many meters are there in 345.6 decimeters?
There are 10 decimetres in one metre. Therefore, 345.6 decimetres is equal to 345.6/10 = 34.56 metres.
What is the largest four-digit number with different digits each greater than the one before it?
6789
What is the mortality rate of gastritis in southeast asian?
The question cannot be answered since it ends in an adjective without the defining noun. Southeast Asian what: SE Asian men or SE Asian women or SE Asian children or SE Asian cats?
Where is the marble in Bugdon 2?
A lot of people have a difficult time getting over the red walls. This seems to mostly affect people on slower computers since timing is critical, and at slow frame rates it is hard to do the jump. The best way to get over these walls is to simply jump on top of one of the wooden blocks. Then jump again and start flying at the apex of the jump. You should now be able to just flyover the red wall.Many people have a hard time finding the marble. The passage to the marble is narrow, but there is a line of acorns marking the entrance. Just keep going to the end of the passage and the marble will be there.The correct way to do the marble bowling is to drop the marble and kick it into the batteries, however, you can also get a running start and just drop the marble. That will send it coasting into the batteries too, and this method is often more accurate than kicking.The slot car racing is really a no-brainer. Just hold down the throttle until the race is over - you'll win every time.
What is the probability of drawing 2 similar cards from a deck of 5?
The answer depends on how many of the 5 cards are similar.
How do you find the outlier in a Stem and Leaf Plots problem?
Usually it would be on a stem all by itself, a long way from the next stem.
What is the probability of rolling a 4 or a 12 on the first roll of a pair of dice?
The probability of rolling a 4 is 3/36 .
The propbability of rolling 12 is 1/36 .
The probability of rolling either one is 4/36 = 1/9 = 111/9 percent.
What does numerical data vary mean?
When there are a variety of different numbers numerical data varies.
-RJC
We have to use the expression P(X=x) = nCx px (1--p)(n--x)
Here n = n and p=p and x = 1 or x>1
P(X>/=1) = 1 -- P(X</=1)
So, P(X<=1) = P(X=0) + P(X=1)
This gives nC0 p0 (1--p)(n-0)+ nC1 p1 (1--p)(n--1)
ie (1--p)n + n p (1--p)(n--1)
A type of selection where the normal distribution is pushed apart into two separate peaks?
Bimodal Distribution
What is the probability of picking a 5 in a dice?
The probability of any one number on a die being rolled is 1/6 or 16.67%.
Identify 4 possible outcomes of a bio logical agent attack?
contaminate equipment
overwhelm med staff and resources
force personnel to use protective equip that degrades mission effectiveness
incapacitate and /or kill large number of personnel